Antacids are drugs that are used to combat hyperacidity by neutralizing the excess acid (HCl) present in the stomach. Another popular component of antacid preparations is aluminum hydroxide, Al(OH)3.

If the neutralization reaction is only 94.3% efficient, how many grams of HCl can be actually neutralized by one tablespoon of the preparation?
1
Expert's answer
2020-11-23T07:32:51-0500

Al(OH3) +3HCl = AlCl3 + 3H2O

1 mole of Al(OH3) = 3 moles of HCl

78g of Al(OH3) = 3 × 36.5g HCl

1g of Al(OH3) = 3 × 36.5 / 78 g of HCl

1 tablespoon = 25g

1 tablespoon of Al(OH3) = 3 ×36.5 × 25 / 78 g of HCl

= 35.1g HCl

This is the theoretical yield of HCl


% efficiency = actual yield/theorectical yield ×100%

94.3% = a / 35.1 ×100

a = 94.3% × 35.1/ 100% = 33.10g of HCl